# references.bib

### v1

@@ IV-B Exact Recovery under Sign-restricted Disturbances
$\hskip 85.35826pt\leq\exp\biggr{(}\lambda^{2}T\Bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\Bigr{)}^{2}\biggr{)}$ (23b)
for all $\lambda\in\mathbb{R}$, where $\bar{C}=6\sqrt{e}\Bigr{(}1+\sqrt{\frac{\pi}{\log 2}}\Bigr{)}$.
Proof:
Recall the definition (7) of the $\psi_{2}$-norm of sub-Gaussian vectors. For a matrix $A\in\mathbb{R}^{d\times d}$ and a vector $x\in\mathbb{R}^{d}$, we have
$\|Ax\|_{\psi_{2}}$ $=\sup_{\|y\|_{2}\leq 1,y\in\mathbb{R}^{d}}\|x^{T}A^{T}y\|_{\psi_{2}}$
$\leq\sup_{\|A^{T}y\|_{2}\leq\|A\|_{2},y\in\mathbb{R}^{d}}\|x^{T}A^{T}y\|_{\psi_{2}}$
$\leq\sup_{\|z\|_{2}\leq\|A\|_{2},z\in\mathbb{R}^{d}}\|x^{T}z\|_{\psi_{2}}$
$=\sup_{\|\tilde{z}\|_{2}\leq 1,\tilde{z}\in\mathbb{R}^{d}}\Bigr{\|}\|A\|_{2}\cdot x^{T}\tilde{z}\Bigr{\|}_{\psi_{2}}=\|A\|_{2}\cdot\|x\|_{\psi_{2}},$ (24)
where the first inequality comes from the relationship
and the last equality is by the homogeneity of the $\psi_{2}$-norm.
Now, due to the system dynamics (9), we have
$x_{t}=(A^{*})^{t}x_{0}+\sum_{i=0}^{t-1}(A^{*})^{t-i-1}w_{i},$ (25)
for $t=0,1,\dots,T$. Then, we obtain
$\|x_{t}\|_{\psi_{2}}$ $=\Bigr{\|}(A^{*})^{t}x_{0}+\sum_{i=0}^{t-1}(A^{*})^{t-i-1}w_{i}\Bigr{\|}_{\psi_{2}}$
$\leq\|(A^{*})^{t}x_{0}\|_{\psi_{2}}+\sum_{i=0}^{t-1}\|(A^{*})^{t-i-1}w_{i}\|_{\psi_{2}}$
$\leq\|(A^{*})^{t}\|_{2}\cdot\|x_{0}\|_{\psi_{2}}+\sum_{i=0}^{t-1}\|(A^{*})^{t-i-1}\|_{2}\cdot\|w_{i}\|_{\psi_{2}}$
$\leq\|A^{*}\|_{2}^{t}\cdot\|x_{0}\|_{\psi_{2}}+\sum_{i=0}^{t-1}\|A^{*}\|_{2}^{t-i-1}\cdot\|w_{i}\|_{\psi_{2}}$
$\leq\sigma_{w}\sum_{i=0}^{t}\|A^{*}\|_{2}^{i}\leq\frac{\sigma_{w}}{1-\|A^{*}\|_{2}},$
where the first two inequalities are due to the triangle inequality and (24), respectively, and the last two inequalities come from Assumptions 2 and 1, respectively. Considering standard basis vectors in $\mathbb{R}^{d}$, the definition (7) states that the $\psi_{2}$-norms of one-dimensional projections satisfy $\|x_{t}^{j}\|_{\psi_{2}}\leq\|x_{t}\|_{\psi_{2}}\leq\frac{\sigma_{w}}{1-\|A^{*}\|_{2}}$ for $j=1,\dots,d.$ Since the definition (2) involves the squared variable, alternating the sign or taking the absolute value does not affect the norm of sub-Gaussian variables. Thus, given the filtration $\mathcal{F}_{t}$, we have
$\|z_{t}^{ij+}\|_{\psi_{2}}\leq\frac{\sigma_{w}}{1-\|A^{*}\|_{2}},\quad\|z_{t}^{ij-}\|_{\psi_{2}}\leq\frac{\sigma_{w}}{1-\|A^{*}\|_{2}}$ (26)
conditioned on $\mathcal{F}_{t}$ for all $i,j\in\{1,\dots,d\}$ and all $t$. Using Lemma 1 in turn, one arrives at
$\|z_{t}^{ij+}-\mathbb{E}[z_{t}^{ij+}]\|_{\psi_{2}}\leq\Bigr{(}1+\sqrt{\frac{\pi}{\log 2}}\Bigr{)}\frac{\sigma_{w}}{1-\|A^{*}\|_{2}}$ (27)
conditioned on $\mathcal{F}_{t}$ for all $i,j\in\{1,\dots,d\}$ and all $t$. The same inequality also holds for $z_{t}^{ij-}$. Notice that the expectation of $z_{t}^{ij+}-\mathbb{E}[z_{t}^{ij+}]$ is zero due to the tower rule and we can use the equivalent Definition 2 to obtain
@@
where $\bar{C}=6\sqrt{e}\Bigr{(}1+\sqrt{\frac{\pi}{\log 2}}\Bigr{)}$. Then, we have
$\mathbb{E}\biggr{[}\exp\Bigr{(}\lambda\Bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}\Bigr{]}\Bigr{)}\biggr{]}$
$=\mathbb{E}\biggr{[}\mathbb{E}\Bigr{[}\exp\Bigr{(}\lambda\Bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}\Bigr{]}\Bigr{)}~{}\Bigr{|}~{}\mathcal{F}_{T-1}\Bigr{]}\biggr{]}$
$=\mathbb{E}\biggr{[}\Bigr{(}\sum_{t=0}^{T-2}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-2}z_{t}^{ij+}\bigr{]}\Bigr{)}$
$\hskip 56.9055pt\cdot\mathbb{E}\Bigr{[}\exp\Bigr{(}\lambda\Bigr{[}z_{T-1}^{ij+}-\mathbb{E}\bigr{[}z_{T-1}^{ij+}\bigr{]}\Bigr{]}\Bigr{)}~{}\Bigr{|}~{}\mathcal{F}_{T-1}\Bigr{]}\biggr{]}$
$\leq\mathbb{E}\biggr{[}\Bigr{(}\sum_{t=0}^{T-2}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-2}z_{t}^{ij+}\bigr{]}\Bigr{)}\biggr{]}\exp\biggr{(}\lambda^{2}\Bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\Bigr{)}^{2}\biggr{)}$
$\leq\dots\leq\exp\biggr{(}\lambda^{2}T\Bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\Bigr{)}^{2}\biggr{)}$
for all $i,j\in\{1,\dots,d\}$, by repetitively applying the tower rule ($z_{t-1}^{ij+}$ is a function of $\mathcal{F}_{t}$) and (28). Due to (26), the same inequality also holds even when $z_{t}^{ij+}$ is replaced by $z_{t}^{ij-}$. ∎
Now, we present one of our main theorems built on Lemmas 3 and 4.
Theorem 2
Suppose that Assumptions 1, 2, 3, 4, and 5 hold. Given $\delta\in(0,1],$ the $l_{1}$-norm estimator defined in (11) achieves the exact recovery after
$T\geq O\biggr{(}\biggr{(}\frac{\sigma_{w}}{(1-p)\lambda(1-\|A^{*}\|_{2})}\biggr{)}^{2}\biggr{)}\cdot\log\Bigr{(}\frac{2d^{2}}{\delta}\Bigr{)}$ (29)
with probability at least $1-\delta$.
Proof:
By Lemma 4, $\sum_{t=0}^{T-1}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}$ satisfies (3) with the parameter $\sqrt{T}\bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\bigr{)}$ for all $i,j\in\{1,\dots,d\}$. Then, (6) in Lemma 2 implies that for all $s>0$, we have
$\mathbb{P}\Bigr{(}\sum_{t=0}^{T-1}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}>-s\Bigr{)}$
$\hskip 85.35826pt\geq 1-\exp\biggr{(}-\frac{s^{2}}{4T\bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\bigr{)}^{2}}\biggr{)}.$
Since the positive expectation is guaranteed by Lemma 3, one can substitute $\frac{\mathbb{E}[\sum_{t=0}^{T-1}z_{t}^{ij+}]}{2}>0$ into $s$ to obtain
$\mathbb{P}\Bigr{(}\sum_{t=0}^{T-1}z_{t}^{ij+}\geq 0\Bigr{)}$ $\geq\mathbb{P}\Bigr{(}\sum_{t=0}^{T-1}z_{t}^{ij+}>\frac{\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}}{2}\Bigr{)}$
$\geq 1-\exp\biggr{(}-\frac{\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}^{2}}{16T\bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\bigr{)}^{2}}\biggr{)}$
$\geq 1-\exp\biggr{(}-\frac{((1-p)\lambda T)^{2}}{16T\bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\bigr{)}^{2}}\biggr{)}$ (30)
for all $i,j\in\{1,\dots,d\}.$ The same lower bound holds for $\sum_{t=0}^{T-1}z_{t}^{ij-}$. In light of the union bound, both (14a) and (14b) hold for all $i,j\in\{1,\dots,d\}$ with probability at least $1-\delta$ if (30) is lower-bounded by $1-\frac{\delta}{2d^{2}}$. This provides the sufficient conditions for the exact recovery with probability at least $1-\delta$ due to Theorem 1. Thus, it suffices to require
$\hskip 28.45274pt1-\exp\biggr{(}-\frac{((1-p)\lambda T)^{2}}{16T\bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\bigr{)}^{2}}\biggr{)}\geq 1-\frac{\delta}{2d^{2}}$
$\Longleftrightarrow~{}\frac{((1-p)\lambda T)^{2}}{16T\bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\bigr{)}^{2}}\geq\log\Bigr{(}\frac{2d^{2}}{\delta}\Bigr{)}$
$\Longleftrightarrow~{}T\geq\biggr{(}\frac{4\bar{C}\sigma_{w}}{(1-p)\lambda(1-\|A^{*}\|_{2})}\biggr{)}^{2}\cdot\log\Bigr{(}\frac{2d^{2}}{\delta}\Bigr{)}.$
This completes the proof. ∎
Remark 2 (Uniqueness)
In general, multiple solutions can be attained from the minimization problem (11). In such a case, $A^{*}$ may not be the unique solution even after a sufficient time. The uniqueness is violated if the optimality condition is satisfied for some matrix other than $A^{*}$, which can be written as $A^{*}+\tilde{A}$ for some $\mathbb{R}^{d\times d}$. From (15), the solution $A^{*}$ is not unique if
$0\in\sum_{t=0}^{T-1}\partial\|w_{t}-\tilde{A}x_{t}\|_{1}\cdot x_{t}^{T}$ (31)
for some $\tilde{A}\neq 0$. Note that (16) and (31) are the same conditions if all $x_{t}$’s are zero, but this is already prevented by adopting Assumption 4. Furthermore, if the state trajectory $x_{0},\dots,x_{T-1}$ sufficiently spans the state space, which is determined by the rank of true matrix $A^{*}$ and the richness of the disturbance data, we can justify the uniqueness of the solution. Similar analyses have been presented in [zhang2024exact].
IV-C Exact Recovery under Arbitrary Disturbances
In the previous subsection, we have discussed the exact recovery under Assumption 5, which requires an equal probability for the sign of the disturbance to be positive or negative. In real-world applications, requiring exactly the same positive and negative probabilities may be challenging. By adopting the proofs used in the previous case, we will now present a more practical situation without considering Assumption 5: if the probability of having an attack at a given time is less than 0.5, it turns out that $l_{1}$-norm estimator overcomes any adversarial noncentral disturbance and the exact recovery is achieved. We formally state the theorem below.
Theorem 3
Suppose that Assumptions 1, 2, and 4 hold. Also, assume that Assumption 3 holds with $p<0.5$. Given $\delta\in(0,1]$, the $l_{1}$-norm estimator defined in (11) achieves the exact recovery after
$T\geq O\biggr{(}\biggr{(}\frac{\sigma_{w}}{(1-2p)\lambda(1-\|A^{*}\|_{2})}\biggr{)}^{2}\biggr{)}\cdot\log\Bigr{(}\frac{2d^{2}}{\delta}\Bigr{)}$ (32)
with probability at least $1-\delta$.
Proof:
In the scenario of $\mathbb{P}(w_{t}^{i}=0~{}|~{}\mathcal{F}_{t})=p<0.5$, for all $i,j\in\{1,\dots,d\}$, we have
$\mathbb{E}\Bigr{[}$ $\sum_{t=0}^{T-1}z_{t}^{ij+}\Bigr{]}$
$=\sum_{t=0}^{T-1}\mathbb{E}\Bigr{[}|x_{t}^{j}|\cdot\mathbb{P}(w_{t}^{i}=0~{}|~{}\mathcal{F}_{t})$
$\hskip 56.9055pt+x_{t}^{j}\cdot[\mathbb{P}(w_{t}^{i}<0~{}|~{}\mathcal{F}_{t})-\mathbb{P}(w_{t}^{i}>0~{}|~{}\mathcal{F}_{t})]\Bigr{]}$
$\geq\sum_{t=0}^{T-1}\biggr{[}(1-p)\mathbb{E}\bigr{[}|x_{t}^{j}|\bigr{]}-p\cdot\mathbb{E}\bigr{[}|x_{t}^{j}|\bigr{]}\biggr{]}\geq(1-2p)\lambda T>0,$
where the first equality is due to (22) in Lemma 3 and the first inequality comes from
$\bigr{|}\mathbb{P}(w_{t}^{i}<0~{}|~{}\mathcal{F}_{t})-\mathbb{P}(w_{t}^{i}>0~{}|~{}\mathcal{F}_{t})\bigr{|}\leq p,$
and the last inequality comes from $p<0.5$. Similarly, $\mathbb{E}[\sum_{t=0}^{T-1}z_{t}^{ij-}]$ has the same lower bound on the expectation. Since the expectations are positive, one can apply the proof of Lemma 4 to arrive at (32). Note that $1-p$ in (29) is simply replaced by $1-2p$. ∎
Remark 3 (Connection between Theorems 2 and 3)
Recall that Assumption 3 assumes $\mathbb{P}(w_{t}^{i}=0~{}|~{}\mathcal{F}_{t})>0$. In Theorem 2, Assumption 5 requires
$\mathbb{P}(w_{t}^{i}<0~{}|~{}\mathcal{F}_{t})=\mathbb{P}(w_{t}^{i}>0~{}|~{}\mathcal{F}_{t})=\frac{1-\mathbb{P}(w_{t}^{i}=0~{}|~{}\mathcal{F}_{t})}{2},$
which is strictly less than 0.5 for all $i=1,\dots,d$ due to Assumption 3. Now, suppose that every negative $w_{t}^{i}$ is infinitesimal. Then, the probability that $w_{t}^{i}$ is adequately far from zero is equivalent to the probability of positive attack which is less than 0.5, and therefore this approximately satisfies the assumptions of Theorem 3. This implies that if nonzero attacks occur with probability less than 0.5, the $l_{1}$-norm estimator can correctly recover the true system even when those attacks are always positive, which represents the worst-case scenario where the system may easily be deceived towards a positive bias. This surprising result shown by Theorem 3 serves as a special case of Theorem 2.

### v5

@@ IV-B Exact Recovery under Sign-restricted Disturbances
$\leq\cdots\leq\exp\biggr{(}\lambda^{2}(1+|\mathcal{K}_{T}|)\cdot\Theta\Bigr{(}\frac{\sigma_{w}}{1-\|A^{*}\|_{2}}\Bigr{)}^{2}\biggr{)}.$ (24)
Since the term of interest has zero mean, we can use the sub-Gaussian parameter given in (24) and apply the inequality (6b) given in Lemma 1 to derive
$\mathbb{P}\biggr{(}\sum_{\begin{subarray}{c}t=0,\\ w_{t}^{i}\neq 0\end{subarray}}^{T-1}y^{T}x_{t}\cdot\text{sgn}(w_{t}^{i})>-\frac{c\lambda^{5}N_{T}}{2\sigma_{w}^{4}}\biggr{)}$
$\hskip 28.45274pt\geq 1-\exp\Biggr{(}-\Theta\biggr{(}\frac{(\frac{\lambda^{5}N_{T}}{\sigma_{w}^{4}})^{2}}{(1+|\mathcal{K}_{T}|)\cdot(\frac{\sigma_{w}}{1-\|A^{*}\|_{2}})^{2}}\biggr{)}\Biggr{)}$
$\hskip 28.45274pt=1-\exp\Biggr{(}-\Theta\biggr{(}\frac{N_{T}^{2}\cdot(1-\|A^{*}\|_{2})^{2}}{(1+|\mathcal{K}_{T}|)\cdot\tau^{10}}\biggr{)}\Biggr{)}$ (25)
where $c$ and $N_{T}$ are the quantities in Lemma 3. Constructing the union bound using Lemma 3 and (25), we obtain
$\mathbb{P}\biggr{(}\sum_{\begin{subarray}{c}t=0,\\ w_{t}^{i}\neq 0\end{subarray}}^{T-1}y^{T}x_{t}\cdot\text{sgn}(w_{t}^{i})+\sum_{\begin{subarray}{c}t=0,\\ w_{t}^{i}=0\end{subarray}}^{T-1}|y^{T}x_{t}|>\frac{c\lambda^{5}N_{T}}{2\sigma_{w}^{4}}\biggr{)}$ (26)
$\hskip 14.22636pt\geq 1-\exp\Biggr{(}-\Theta\biggr{(}\frac{\lambda^{4}N_{T}}{\sigma_{w}^{4}}\biggr{)}\Biggr{)}$
$\hskip 56.9055pt-\exp\Biggr{(}-\Theta\biggr{(}\frac{N_{T}^{2}\cdot(1-\|A^{*}\|_{2})^{2}}{(1+|\mathcal{K}_{T}|)\cdot\tau^{10}}\biggr{)}\Biggr{)}$ (27)
Finally, due to Assumption 3, the probabilistic attack model implies that we have
$N_{T}\geq\Theta(p(1-p)T),\quad 1+|\mathcal{K}_{T}|\leq\Theta(pT)$ (28)
with probability at least $1-\exp(-\Theta(p(1-p)T))$, considering the expectations of each quantity. The union bound (27) is then lower-bounded by $1-\delta$ if we have
$T\geq\Theta\biggr{(}\max\Bigr{\{}\frac{1}{p(1-p)},\frac{\tau^{4}}{p(1-p)},$
$\hskip 71.13188pt\frac{\tau^{10}}{p(1-p)^{2}(1-\|A^{*}\|_{2})^{2}}\Bigr{\}}\cdot\log\Bigr{(}\frac{1}{\delta}\Bigr{)}\biggr{)},$
which is equivalent to (21) since $1-p$ and $1-\|A^{*}\|_{2}$ are less than $1$.
To satisfy the condition (15) not only for a single $(y,i)$ pair but for all $y\in\mathbb{S}^{d-1}$ and $i\in\{1,\dots,d\}$, we leverage the following lemma presented in [vershynin2010intro].
Lemma 5 (Covering number of the sphere).
For $\epsilon>0$, consider a subset $\mathcal{N}_{\epsilon}$ of $\mathbb{S}^{d-1}$, such that for all $y\in\mathbb{S}^{d-1}$, there exists some point $\tilde{y}\in\mathcal{N}_{\epsilon}$ satisfying $\|y-\tilde{y}\|_{2}\leq\epsilon$. The minimal cardinality of such a subset is called the covering number of the sphere and is upper-bounded by $(1+\frac{2}{\epsilon})^{d}$.
Now, we present one of our main theorems to satisfy sufficient conditions for the exact recovery.
Theorem 2.
Suppose that Assumptions 1, 2, 3, 4, and 5 hold. Given $\delta\in(0,1],$ the $l_{1}$-norm estimator defined in (11) achieves the exact recovery with probability at least $1-\delta$ when
$T\geq\Theta\biggr{(}\frac{d\cdot\max\{1,\tau^{10}\}}{p(1-p)^{2}(1-\|A^{*}\|_{2})^{2}}$
×$\hskip 36.98857pt\times\max\biggr{\{}1,\log\biggr{(}\frac{d\cdot\tau}{\delta(1-p)(1-\|A^{*}\|_{2})}\biggr{)}\biggr{\}}\biggr{)}.$ (29)
Proof.
We first show that the condition (15) holds with high probability, given $i\in\{1,\dots,d\}$. Let
$\epsilon:=\Theta\biggr{(}\frac{(1-p)(1-\|A^{*}\|_{2})}{\tau^{5}\sqrt{\log(1/\delta)}}\biggr{)}.$
For any $y,y^{\prime}\in\mathbb{S}^{d-1}$ such that $\|y-y^{\prime}\|_{2}\leq\epsilon$, we have
$\sum_{t=0}^{T-1}z_{t}^{i}$ $(y)-\sum_{t=0}^{T-1}z_{t}^{i}(y^{\prime})\geq-\sum_{t=0}^{T-1}|(y-y^{\prime})^{T}x_{t}|$
$\geq-\|y-y^{\prime}\|_{2}\sum_{t=0}^{T-1}\|x_{t}\|_{2}\geq-\epsilon\sum_{t=0}^{T-1}\|x_{t}\|_{2}$
$\geq-\epsilon\sum_{t=0}^{T-1}\Bigr{[}\|A^{*}\|_{2}^{t}\cdot\|x_{0}\|_{2}+\sum_{k\in\mathcal{K}_{T}}\|A^{*}\|_{2}^{t-1-k}\|w_{k}\|_{2}\Bigr{]}$
$\geq-\frac{\epsilon}{1-\|A^{*}\|_{2}}\Bigr{[}\|x_{0}\|_{2}+\sum_{k\in\mathcal{K}_{T}}\|w_{k}\|_{2}\Bigr{]}$ (30)
where the first inequality is due to the triangle inequality. Due to homogeneity and the triangle inequality, the sub-Gaussian norm (see (2)) of the term in (30) is bounded by $\frac{\epsilon(1+|\mathcal{K}_{T}|)\sigma_{w}}{1-\|A^{*}\|_{2}}$. We leverage the inequality (5) in Lemma 1 to obtain
$\mathbb{P}\Bigr{(}\sum_{t=0}^{T-1}z_{t}^{i}(y)-\sum_{t=0}^{T-1}z_{t}^{i}(y^{\prime})>-\frac{c\lambda^{5}N_{T}}{4\sigma_{w}^{4}}\Bigr{)}$
$\geq\mathbb{P}\Bigr{(}-\frac{\epsilon}{1-\|A^{*}\|_{2}}\Bigr{[}\|x_{0}\|_{2}+\sum_{k\in\mathcal{K}_{T}}\|w_{k}\|_{2}\Bigr{]}>-\frac{c\lambda^{5}N_{T}}{4\sigma_{w}^{4}}\Bigr{)}$
$\geq 1-2\exp\biggr{(}-\Theta\Bigr{(}\frac{\lambda^{10}N_{T}^{2}(1-\|A^{*}\|_{2})^{2}}{\epsilon^{2}(1+|\mathcal{K}_{T}|)^{2}\sigma_{w}^{10})}\Bigr{)}\biggr{)}$
$=1-2\exp\biggr{(}-\Theta\Bigr{(}\frac{(1-p)^{2}(1-\|A^{*}\|_{2})^{2}}{\epsilon^{2}\tau^{10}}\Bigr{)}\biggr{)}\geq 1-\frac{\delta}{2},$
where the equality is by considering (28), provided that $T\geq\Theta(\frac{1}{p(1-p)}\log(\frac{1}{\delta}))$. Then, by Lemma 5, if we have a subset with $(1+\frac{2}{\epsilon})^{d}$ number of points $\{y_{j}:j=1,2,\dots\}$ in $\mathbb{S}^{d-1}$ that satisfy $\sum_{t=0}^{T-1}z_{t}^{i}(y_{j})>\frac{c\lambda^{5}N_{T}}{2\sigma_{w}^{4}}$ (see (26)) for all $j$ with probability at least $1-\frac{\delta}{2}$, then we will have
$\sum_{t=0}^{T-1}z_{t}^{i}(y)>\frac{c\lambda^{5}N_{T}}{4\sigma_{w}^{4}}>0,\quad\forall y\in\mathbb{S}^{d-1}$ (31)
with probability at least $1-\delta$. Thus, it suffices to replace $\delta$ in (21) with $\frac{\delta}{2(1+2/\epsilon)^{d}}$ to achieve (31). We have
$\log\biggr{(}\frac{2(1+2/\epsilon)^{d}}{\delta}\biggr{)}=\Theta\biggr{(}d\log\biggr{(}1+\frac{2}{\epsilon}\biggr{)}+\log\biggr{(}\frac{1}{\delta}\biggr{)}\biggr{)}$
$=\Theta\biggr{(}d\log\biggr{(}1+\frac{2\tau^{5}\sqrt{\log(1/\delta)}}{(1-p)(1-\|A^{*}\|_{2})}\biggr{)}+\log\biggr{(}\frac{1}{\delta}\biggr{)}\biggr{)}$
$\leq\Theta\biggr{(}\max\biggr{\{}d+\log\biggr{(}\frac{1}{\delta}\biggr{)},d\log\biggr{(}\frac{2\tau^{5}\sqrt{\log(1/\delta)}}{\delta(1-p)(1-\|A^{*}\|_{2})}\biggr{)}\biggr{\}}\biggr{)}$
$=\Theta\biggr{(}\max\biggr{\{}d,d\log\biggr{(}\frac{\tau}{\delta(1-p)(1-\|A^{*}\|_{2})}\biggr{)}\biggr{\}}\biggr{)},$ (32)
The last equality is because the first case in the max argument arises when
$\frac{2\tau^{5}\sqrt{\log(1/\delta)}}{(1-p)(1-\|A^{*}\|_{2})}\leq 1,$
which implies $\log(\frac{1}{\delta})\leq\Theta(\frac{1}{\tau^{10}})$ since $1-p$ and $1-\|A^{*}\|_{2}$ are less than 1. Note that Assumptions 2 and 4 result in
$\lambda^{2}d$ $\leq\text{tr}(\mathbb{E}[w_{t}w_{t}^{T}~{}|~{}\mathcal{F}_{t}])=\mathbb{E}[\text{tr}(w_{t}w_{t}^{T})~{}|~{}\mathcal{F}_{t}]$
$=\mathbb{E}[\|w_{t}\|^{2}_{2}~{}|~{}\mathcal{F}_{t}]\leq 18\bigr{\|}\|w_{t}\|_{2}\bigr{\|}_{\psi_{2}}\leq 18\bigr{\|}w_{t}\bigr{\|}_{\psi_{2}}\leq 18\sigma_{w}^{2}$
conditioned on any $\mathcal{F}_{t}$, where the second inequality is due to (3) and the third inequality comes from Definition 3. Thus, $\frac{1}{\tau^{10}}$ is bounded by the quantity $(\frac{18}{d})^{5}$, which yields $d+\log(\frac{1}{\delta})=\Theta(d)$. Furthermore, the second case in the max argument arises from $\frac{1}{\delta}$ dominating the term $\sqrt{\log(1/\delta)}$. To conclude, we achieve (31) with probability at least $1-\delta$ after
$T\geq\Theta\biggr{(}\frac{d\cdot\max\{1,\tau^{10}\}}{p(1-p)^{2}(1-\|A^{*}\|_{2})^{2}}$
×$\hskip 36.98857pt\times\max\biggr{\{}1,\log\biggr{(}\frac{\tau}{\delta(1-p)(1-\|A^{*}\|_{2})}\biggr{)}\biggr{\}}\biggr{)}.$ (33)
Letting (31) (or equivalently, (15)) hold with probability at least $1-\frac{\delta}{d}$, and thus replacing $\delta$ in (33) with $\frac{\delta}{d}$, suffices to guarantee that the condition (15) holds for all $i\in\{1,\dots,d\}$ with probability at least $1-\delta$. By Theorem 1, the exact recovery is achieved with the desired probability.
Remark 2 (Role of attacks):
In Theorem 2, the required time for exact recovery is $\tilde{\Theta}\bigr{(}\frac{d}{p(1-p)^{2}}\bigr{)}$, which is minimized when the attack probability $p$ is 1/3. This stems from the role of attacks in two orthogonal ways. First, the estimator requires a longer time as $p\to 1$ since the attacker tries to deceive the system. However, since the attack is also a source of excitement due to Assumption 4, the estimator finds it difficult to learn the system through the attacks as $p\to 0$. For example, if $x_{0},w_{0},\dots,w_{T-1}$ are all identically zero, any matrix $A$ serves as a solution to the estimator, but our goal is to guarantee that the true matrix is the unique optimizer.
IV-C Exact Recovery under Arbitrary Disturbances
In the previous subsection, we have discussed the exact recovery under Assumption 5, which requires an equal probability for the sign of the disturbance. In real-world applications, requiring exactly the same positive and negative probabilities may be challenging. We will now present a more practical situation without considering Assumption 5: if the attack probability is less than 0.5, it turns out that the $l_{1}$-norm estimator overcomes arbitrary (possibly adversarial) noncentral disturbance and the exact recovery is achieved. This includes the case where the arbitrary attacks are always positive (or negative), representing the worst-case scenario in which the system may easily be deceived towards a unidirectional bias. We formally state the theorem below.
Theorem 3.
Suppose that Assumptions 1, 2, and 4 hold, and Assumption 3 holds with $0.50<p<$ 0.5. Given $\delta\in(0,1]$, the $l_{1}$-norm estimator defined in (11) achieves the exact recovery with probability at least $1-\delta$ when
$T\geq\Theta\biggr{(}\frac{d\cdot\max\{1,\tau^{10}\}}{p(1-2p)^{2}(1-\|A^{*}\|_{2})^{2}}$
×$\hskip 31.29802pt\times\max\biggr{\{}1,\log\biggr{(}\frac{d\cdot\tau}{\delta(1-2p)(1-\|A^{*}\|_{2})}\biggr{)}\biggr{\}}\biggr{)}.$ (34)
Proof.
We first focus on a restricted class of problems where the distribution of $w_{t}$ can be constructed via two types of attack vectors $v_{t}$ and $\tilde{v}_{t}$ as:
$w_{t}\sim I_{A}\cdot[I_{B}\cdot v_{t}+(1-I_{B})\cdot\tilde{v}_{t}],$ (35)
$I_{A}\sim\text{Bernoulli}(2p),\quad I_{B}\sim\text{Bernoulli}(0.5),$
$I_{A}\perp\!\!\!\perp I_{B},\quad(I_{A},I_{B})\perp\!\!\!\perp v_{t},\quad(I_{A},I_{B})\perp\!\!\!\perp\tilde{v}_{t}$
@@ IV-C Exact Recovery under Arbitrary Disturbances
In this scenario, Assumptions 3, 4, and 5 are satisfied with the aggregate attack probability $10<2p<$ 1. Under Assumptions 1 and 2, one can now apply Theorem 2 to obtain (3), where $1-p$ in (2) is replaced with $1-2p$. This results from the modification of (28) as
$N_{T}\geq\Theta(p(1-2p)T),\quad 1+|\mathcal{K}_{T}|\leq\Theta(2pT)$ (36)
with high probability, since the lower bound of $N_{T}$ defined in Lemma 3 accounts for the product of the probabilities of no attack ($1-2p$) and an attack satisfying Assumption 4 (in this case, $v_{t}$ occurs with probability $p$). The upper bound of $1+|\mathcal{K}_{T}|$ should consider the total attack probability $2p$.
Now, suppose that $\tilde{v}_{t}$ in (35) is designed as $\tilde{\alpha}_{t}\circ\tilde{\beta}_{t}$, with $\tilde{\alpha}_{t}$ having finite entries. Then, let $\tilde{\beta}_{t}\to 0^{+}$, meaning that all coordinates $\tilde{\beta}_{t}^{i}$ can be arbitrarily small but remain positive, and thus $\tilde{v}_{t}^{i}$ will maintain their signs with arbitrarily small magnitudes. This ensures that Assumption 5 is not violated and (31) still holds after the recovery time (3). Taking the limit $\tilde{\beta}_{t}\to 0^{+}$ can be extended to the case where $\tilde{\beta}_{t}=0$, at which point the distribution (35) collapses to $w_{t}\sim I_{A}\cdot I_{B}\cdot v_{t}$, with $0.50<p<$ 0.5 and $v_{t}$ satisfying Assumption 4. This is consistent with the assumptions for $w_{t}$ given by this theorem, where the inequality (31) is revised to a non-strict form as
$\sum_{t=0}^{T-1}z_{t}^{i}(y)=\sum_{\begin{subarray}{c}t=0,\\ w_{t}^{i}\neq 0\end{subarray}}^{T-1}(y^{T}x_{t}\cdot\text{sgn}(w_{t}^{i}))+\sum_{\begin{subarray}{c}t=0,\\ w_{t}^{i}=0\end{subarray}}^{T-1}|y^{T}x_{t}|$
$\hskip 91.04881pt\geq\frac{c\lambda^{5}N_{T}}{4\sigma_{w}^{4}}>0,\quad\forall y\in\mathbb{S}^{d-1},$ (37)
since $y^{T}x_{t}$ is a continuous function of $w_{0},\dots,w_{t-1}$ and $y^{T}x_{t}\cdot\text{sgn}(w_{t}^{i})$ either increases to or remains at $|y^{T}x_{t}|$ when $\tilde{\beta}_{t}^{i}\to 0^{+}$ is converted to $\tilde{\beta}_{t}^{i}=0$ (see (14)). Note that the last inequality of (IV-C) remains strict, ensuring that the condition (15) is satisfied with the exact recovery time (3).
Remark 3:
In Theorem 3, we highlight that the $l_{1}$-norm estimator withstands arbitrary adversarial noncentral disturbances with the exact recovery time of $\tilde{\Theta}\bigr{(}\frac{d}{p(1-2p)^{2}}\bigr{)}$, provided that the attack probability is restricted to $0.50<p<$ 0.5.